OpenStudy (anonymous):

solve the proportion x/2=5/15

hartnn (hartnn):

hint : 5 goes into 15 15 = 5*3

hartnn (hartnn):

so what will be 5/15 = .. ?

OpenStudy (anonymous):

\[\frac{ x }{ 2 } = \frac{ 5 }{ 15 } =>x = \frac{ 5 }{ 15 } \times 2 =\frac{ 2 }{ 3 }\] x = 2/3